Solve the equation using factoring techniques.
10x^2-19x+6=0.

Found 2 solutions by Nate, jake_6233:
Answer by Nate(3500)

10x^2 - 19x + 6 = 0 factors of 60 that sum to make -19 .... factors: -4 and -15
10x^2 - 4x - 15x + 6 = 0
(10x^2 - 4x) + (-15x + 6) = 0
2x(5x - 2) + -3(5x - 2) = 0
(2x - 3)(5x - 2) = 0
2x - 3 = 0 and 5x - 2 = 0
2x = 3 and 5x = 2
x = 1.5 and x = 0.4

Answer by jake_6233(74)

10x^2-19x+6=0
(2x-3)(5x-2)=0
2x-3=0 x=3/2
5x-2=0 x=2/5
You may not have seen this factoring technique but you know the rule for factoring if the coeficcient of x^2 is one right? Well look for two numbers that multiply to 60 and add to -19 We have (10x-15)(10x-4) (These won't give you the polynomial in your problem yet) Factor out a 5 from (10x-15) and a 2 from (10x-4) Then you have (2x-3)(5x-2) It is a bit quicker than trial and error